OpenStudy (anonymous):

Choose the graph which matches the following system of equations. x + y = 4 2x + 3y = 18

OpenStudy (anonymous):

hmm what graph?

OpenStudy (anonymous):

Shouldn't you solve this system first..

OpenStudy (anonymous):

x + y = 4 2x + 3y = 18 x=4-y 2(4-y)+3y=18 8-2y+3y=18 y=10 x=4-y x=4-10 x=-6

OpenStudy (anonymous):

so how do i know which graph presents that information? I'm learning this.

OpenStudy (anonymous):

Do you know how to plot?

OpenStudy (anonymous):

y=10 x=-6 Find these points on a graph, draw a line connecting them

OpenStudy (anonymous):

ohhhh okay got it!!!! thanks!